OpenStudy (anonymous):

Find cot θ if sec θ =2.

OpenStudy (anonymous):

@Luigi0210 Can you help me with this last one please?

OpenStudy (psymon):

Do you know what sides of a triangle secant corresponds to?

OpenStudy (luigi0210):

Yup, draw a triangle

OpenStudy (anonymous):

I know secant is H/A .

OpenStudy (anonymous):



OpenStudy (luigi0210):

Woah, so many people answering..

OpenStudy (anonymous):

Unexpectedly, yes lol.

OpenStudy (psymon):

Right. So if H/A = 2, then A must be 1. with that knowledge, use those two values and pythagorean theorem to find thefinal side. Then do you know what sides cotangent are?

OpenStudy (anonymous):

Cot is A/O

OpenStudy (luigi0210):

Well I guess I should back off..

OpenStudy (mertsj):



OpenStudy (mertsj):

cot = side adj/side opp=1/sqrt3=sqrt3/3

OpenStudy (anonymous):

Hold on I'm confused. How do I label the secant θ = 2 first?

OpenStudy (mertsj):

secant = hypotenuse/adjacent=2/1

OpenStudy (mertsj):

Use Pythagorean Theorem to find missing side.

OpenStudy (anonymous):

It would be.. \[\sqrt{2}\]